OpenStudy (anonymous):

Let sin x = .22. Which sec x? A. 1.03 B. 1.13 C. 1.33 D. 1.30

OpenStudy (anonymous):

You should really give it a try yourself first..... Sove for x arcsin0.22 Do I/cosx... you will get the answer.

OpenStudy (anonymous):

I think arcsine(0.22) = angles 12.709 and 167.29. So if the value of x = 12.709. So sec x= 1/cosx ... could you explain your train of thought? @Jhannybean

OpenStudy (jhannybean):

I see where you were going with that.

OpenStudy (anonymous):

I thought the answer was 1.03..... am I wrong?

OpenStudy (jhannybean):

Nevermind :) The question looked a little confusing to me. So we'll do \[\large \sin^{-1}(0.22)=x \] We get our x-value. Using this x-value, we plug it into 1/cos(x) found by the identity \(\large \sec(x) = \cfrac{1}{\cos(x)}\)plugging in the x-value we will have \[\large \cfrac{1}{\cos(\sin^{-1}(0.22))} =?\]

OpenStudy (anonymous):

Few I though I was the one that was confused ;) my skills haven't left me yet haha

OpenStudy (jhannybean):

\[\large \sin^2(x) +\cos^2(x) = 1\]\[\large \sin(x) = 0.22\]\[\large (0.22)^2 +\cos^2(x) =1\]\[\large \cos^2(x) = 1-(0.22)^2\]\[\large \cos(x) = \pm \sqrt{1-(0.22)^2}\]And since 0.22 is in quad 1,we're only taking positive values, so, \[\large \cos(x) = +\sqrt{1-(0.22)^2}\]We know that \(\large \sec(x) = \cfrac{1}{\cos(x)}\) So apply this to the problem.\[\large \sec(x) = \cfrac{1}{\sqrt{1-(0.22)^2}}\]

OpenStudy (anonymous):

how do you make such nice bold print writing.. Its looks very professional @Jhannybean

OpenStudy (jhannybean):

equation editor down below :)

OpenStudy (anonymous):

but look when I type it looks like this...\[\sin ^{2} +\cos ^{2} = 1\]

OpenStudy (anonymous):

o shoot nvm....